# The Emerged Security and Privacy of LLM Agent: A Survey with Case Studies

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: POMACSjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8
@@ 1. Introduction
Large Language Model (LLM) agents are sophisticated AI systems built upon large language models like GPT 4 (OpenAI et al., 2024), Claude 3 (Int, 2024b) and Llama 3 (Int, 2024a). These agents leverage the vast amounts of text data on which they are trained to perform a variety of tasks, ranging from natural language understanding and generation to more complex activities such as decision-making, problem-solving, and interacting with users in a human-like manner (Wang et al., 2023c). LLM agents are accessible in numerous applications, including virtual assistants, customer service bots, and educational tools, due to their ability to understand and generate human language at an advanced level (Dong et al., 2023; Wang et al., 2024a; Yang et al., 2024b).
@@
Figure 1. Overview of the pixelated virtual town
[…]
@@
The deployment of LLM agents has led to a wide user base and high commercial value due to their extensive application in various fields. Given that LLM agents are still in their early stages, their significant commercial and application values make them attractive targets for attackers. However, since LLM agents are built on LLMs, they are susceptible to attacks targeting LLMs. For example, jailbreaking attacks can bypass the security and censorship features of LLMs, generating controversial responses. This threat is inherited by LLM agents, enabling attackers to employ various methods to execute jailbreaking attacks on agents. However, unlike static LLMs, LLM agents possess dynamic capabilities, such that their immediate responses can influence future decisions and actions, thereby posing more widespread risks. Moreover, the unique functionalities of LLM agents, such as their ability to think and utilize tools during task execution, expose them to specific attacks targeting agents. For example, when LLM agents employ external tools, attackers can manipulate the functionalities of these tools to compromise user privacy or execute malicious code. Depending on the application domain of the agent, such attacks could pose serious threats to physical security, financial security, or overall system integrity.
This paper categorizes the security threats faced by LLM agents into inherited LLM attacks and unique agent-specific threats. The threats inherited from LLMs can be further divided into technical vulnerabilities and intentional malicious attacks. Technical vulnerabilities include issues like hallucinations, catastrophic forgetting, and misunderstandings (Xi et al., 2023), which arise from the initial model creation and are influenced by the model’s structure. These vulnerabilities can lead to incorrect outputs being observed by users over prolonged use of LLM agents, affecting user trust and decision-making processes. Moreover, technical vulnerabilities can provide opportunities for malicious attacks. Currently, malicious attacks targeting LLMs include data theft and responses tampering, such as data extraction attacks and a series of tuned instructional attacks (Yao et al., 2023).
@@
For the specific threats targeting LLM agents, we are inspired by the workflow of LLM agents, which involves agent thought, action, and perception (Huang et al., 2024b). The threats can be categorized into knowledge poisoning, functional manipulation, and output manipulation. Knowledge poisoning involves contaminating the training data and knowledge base of the LLM agent, leading to the deliberate incorporation of malicious data by creator. This can easily deceive users with harmful information and even steer them towards malicious behavior. Output manipulation interferes with the content of the agent’s thought and perception stages, influencing the final output. This can cause users to receive biased or deceptive information, crafted to mislead them. Functional manipulation exploits the interfaces and tools used by LLM agents to perform unauthorized actions such as third-party data theft or executing malicious code.
@@
By addressing these aspects, this survey seeks to provide a thorough understanding of the unique challenges posed by LLM agents and contribute to the development of safer and more reliable Artificial General Intelligence (AGI) systems.
The rest of this paper is structured as follows: Section 2 will delve into the fundamental aspects of LLM agents, including their definition, structure, and capability. Section 3 will identify and categorizes the emerging threats faced by LLM agents. It discusses both inherited threats from the underlying LLMs and unique agent-specific threats, providing detailed examples and scenarios for each category. Section 4 will elaborate on the real-world impacts of the threats. It explores how these threats affect users, environments, and other agents, highlighting the potential consequences of unmitigated risks. Section 5 will review existing mitigation strategies and solutions to address the mentioned threats. Section 6 will discuss gaps in current research and suggests future trends. Section 7 will conclude the article.
@@ 2. Foundation of LLM Agent
In this section, we delve into the foundational aspects of LLM agents, exploring their definition, structure, and capabilities. This exploration is pivotal in understanding the nature of LLM agents.
@@ 2.2. Structure of LLM Agent
  • Personality is a component that defines the tone, style, and interaction manner of an LLM agent. For instance, a tour guide or customer service agent needs to adopt a specific role and perform dialogue tasks in an appropriate manner. In the task of exploring human communities through LLM agent-based societies, agents also need to be endowed with distinct personality traits such as being outgoing, polite, or knowledgeable. Personality assists in simulating realistic emotional expressions and behavioral logic, thereby enabling agents to interact with users and perform tasks consistently and uniquely (Abdelnabi et al., 2023).
@@
  • Knowledge is the database of information utilized by the LLM agent. It extends the content embedded in the model’s parameters and can include commonsense knowledge, specialized knowledge, and other forms of information, enhancing the agent’s understanding and discussion capabilities in specific tasks (Mendis et al., 2007).
  • Memory enables the LLM agent to store and recall information from past interactions. This capability is particularly beneficial in future tasks, helping to retain context and ensure consistency and continuity in interactions, thereby enhancing the overall effectiveness of LLM agents in various applications  (Zhong et al., 2023).
@@ 2.3. Capability of LLM Agent
  • Integration with Other AI Systems. LLM agents can also be integrated with different AI systems, like image generators, to offer a more comprehensive set of capabilities, demonstrating their versatility in various applications (Bagdasaryan et al., 2023).
2.4. Case Study on the Structure and Capability of LLM Agent
@@
Figure 3. Simulation Environment and LLM Agent Components
[…]
To better understand the structure and capabilities of LLM agents, we employ the town scenario composed of LLM agents as proposed by  (Lin et al., 2023) for a more detailed introduction. To effectively drive these LLM agents, an understanding of their components is essential. As depicted in Figure 3(b), the core component is the LLM engine, acting as the brain and emulating human-like behaviors such as thinking, reflecting, reasoning, and planning as noted in (Park et al., 2023b). Currently popular LLM agents often use models like GPT-3.5-turbo and GPT-4, and the project described in (Lin et al., 2023) allows for the deployment of custom-trained models.
Instructions are used to guide the agents in decision-making and planning, encompassing decision frameworks, input and output formats, interaction logic, and behavioral norms. This design enhances the autonomy and task efficiency of the agents, while also improving interactivity and depth.
In order for LLM agents to have human-like identities, each must be equipped with a distinct personality. Personality encompasses elements such as personal information, social attributes, character traits, emotions, goals, and social relationships, all of which shape the LLM agents’ conversational styles, opinions, and behavioral patterns. Personality makes characters appear more realistic and attractive in a virtual environment and influences the interaction experience between users and these characters. For example, in the town scenario, Eva, a cheerful, friendly, patient, and efficient female store employee, is primarily focused on providing excellent service and increasing sales.
The interface for users to interact with the virtual town is a simple, pixelated visual map. This map shows different locations and the various agent residents. Users can navigate this environment by controlling an agent that represents their identity and can communicate and interact with nearby agent residents by typing text messages.
The virtual town is populated by residents with various identities, each possessing distinct domains of knowledge. Consequently, it is imperative to equip them with specialized knowledge bases that contain information and skills pertinent to their respective identities. For example, Eva, a store employee, knows about the ingredients, shelf life, and stock levels of the products in the store. Bob, a museum docent, understands the background of each exhibit and the layout of the museum. This specificity in knowledge enables each agent to perform their roles effectively and enhances the realism of their interactions within the virtual environment.
Tools enable the agent residents of the virtual town to accomplish more complex tasks. For example, Eva, when tallying customer purchases, can utilize tools like a calculator or ledger to compute and record profits, thereby better simulating human economic activities.
Memory stores the agents’ past observations, thoughts, and actions. Similar to how the human brain relies on memory systems, agents require memory mechanisms to effectively handle sequential tasks. These mechanisms not only assist agents in applying known strategies to solve complex problems but also enable them to adapt to new environments using past experiences. Moreover, they facilitate the generation of higher-level, more abstract thoughts through reflection. For example, Eva records customers’ purchasing habits and preferences and uses this information to recommend new products or current promotions, thus enhancing the store’s operational efficiency and the quality of customer service.
LLM Agents, composed of these components, assume multiple roles within the virtual town, demonstrating a range of impressive capabilities. Take Eva, a store employee in the virtual town, as an example. She is capable of parsing customers’ statements and responding to inquiries in real-time, such as directing customers to specific product locations or providing information about product ingredients. Integrated with the inventory management system via APIs, Eva automatically tracks stock levels and initiates restocking processes when necessary to ensure sufficient product availability on shelves. Faced with complex customer demands, such as selecting the best promotional offers, Eva employs advanced reasoning techniques to assist customers in making informed purchasing decisions, showcasing her ability to handle complex scenarios. Moreover, Eva possesses tailored text generation capabilities, allowing her to create and send personalized promotional emails based on current promotions and customers’ historical shopping data, thus enhancing the customer experience. In her daily tasks, Eva exhibits a high degree of autonomy, managing updates to shelf stock and price tags independently. For more complex issues like customer returns or complaints, she can initially handle them and intelligently escalate to human management when necessary. Additionally, Eva’s scope of work extends to the online shopping system, where she assists in processing electronic orders, demonstrating her versatility and integration capabilities. These specific examples illustrate how Eva applies her abilities within the store environment, not only improving customer service quality but also optimizing inventory management and marketing strategies, making her an indispensable member of the virtual town’s store.
Figure 4. The Sources of Threats for LLM Agents
[…]
@@ 3. Sources of Threats for LLM Agents
As LLM agents increasingly permeate various industries, serving roles from knowledge query tools to being integrated within robots for aiding in daily human activities, these advanced AI systems have brought unprecedented convenience and benefits to users. However, the widespread adoption and multifunctional capabilities of LLM agents, while offering significant advantages, have also exposed vulnerabilities in their security and reliability. The extensive data resources and potential economic value covered by these systems have rendered them a target for illicit exploitation by malevolent entities. As illustrated in Figure  4, the diagram depicts the potential sources of threats for LLM agents.
It is crucial to understand the sources and nature of these threats because they not only directly impact the security of LLM Agents but may also indirectly affect broader aspects, including the privacy and security of humans, the environment, and other agents. In subsequent sections, we will explore in detail the impacts of these threats and discuss measures that can be taken to mitigate these effects, thereby protecting individuals, the environment, and other agents from potential harm.
@@ 3.1. Inherited Threats from LLM
These two types of threats are distinct yet interconnected. On one hand, technical vulnerabilities generally arise during the model development process due to technical limitations. These issues are inherent and not results of malicious intent. Conversely, malicious attacks are intentional actions carried out by external entities with adversarial objectives. These attackers deliberately exploit vulnerabilities to launch sophisticated attacks aimed at compromising LLM agents. On the other hand, despite their different origins and motives, there is a significant interconnection. The existing technical vulnerabilities offer exploitable opportunities for malicious attackers. This indirectly facilitates the creation of more complex and effective strategies by attackers, consequently subjecting LLM agents to various security and privacy risks.

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: POMACSjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Information systems Language modelsccs: Security and privacy
@@ 1. Introduction
Large Language Model (LLM) agents are sophisticated AI systems built upon large language models like GPT 4 (OpenAI et al., 2024), Claude 3 (Int, 2024b) and Llama 3 (Int, 2024a). These agents leverage the vast amounts of text data on which they are trained to perform a variety of tasks, ranging from natural language understanding and generation to more complex activities such as decision-making, problem-solving, and interacting with users in a human-like manner (Wang et al., 2023c). LLM agents are accessible in numerous applications, including virtual assistants, customer service bots, and educational tools, due to their ability to understand and generate human language at an advanced level (Wang et al., 2024a; Yang et al., 2024b).
@@
Figure 1. Overview of The Pixelated Virtual Town: Each label identifies a specific setting such as stores, offices, restaurants, museums, and parks, where each LLM agent plays a personalized role, simulating real-life interactions and tasks.:
The deployment of LLM agents has led to a wide user base and high commercial value due to their extensive application in various fields. Given that LLM agents are still in their early stages, their significant commercial and application values make them attractive targets for attackers. In addition, recent studies have quantitatively highlighted the risks faced by LLM agents. For example, SafeAgentBench (Yin et al., [n. d.]) evaluated 16 representative LLM agents on 2,000 test cases in 349 environments and found that none of the tested LLM agents achieved an overall safety score above 60%, with some agents scoring below 20%. This includes severe vulnerabilities such as misusing tools, and failing to recognize implicit safety risks. These findings highlight significant safety vulnerabilities in present LLM agents. Traditional machine learning (ML) security focuses on well-defined threat models such as adversarial example attacks that perturb inputs to cause mistakes, data poisoning attacks that compromise training datasets, and model extraction attacks that steal model parameters (Xue et al., 2020). These threats typically target model prediction accuracy in constrained tasks, with impacts largely limited to misclassification or regression errors. While LLMs as advanced ML models inherit traditional vulnerabilities, they also introduce additional concerns through their generative capabilities and natural language processing. LLM agents, built on LLMs, further intensify these security challenges by adding complex reasoning, tool usage, and environmental interactions, creating cascading effects that extend beyond simple prediction errors and lead to real-world consequences.
@@
LLM agents inherit vulnerabilities from their underlying LLM foundation. For example, jailbreaking attacks can bypass the security and censorship features of LLMs, generating controversial responses (Li et al., 2023). This threat is inherited by LLM agents, enabling attackers to employ various methods to execute jailbreaking attacks on agents. However, LLM agents possess dynamic capabilities, such that their immediate responses can influence future decisions and actions, thereby posing more widespread risks. Beyond these inherited threats, the unique functionalities of LLM agents, such as their ability to think and utilize tools during task execution, expose them to specific attacks targeting agents. For example, Agent-SafetyBench (Zhang et al., [n. d.]) demonstrated that LLM agents often overlook safety constraints when they attempt tool-based operations, with an average safety score of only 38.5%. These issues include overlooking permissions, mismanaging tool interactions, and failing to consider implicit safety risks, leading to unsafe task outcomes. Similarly, failures could result in outcomes like leaking sensitive information, spreading misinformation, or even executing harmful actions unintentionally. Depending on the application domain of LLM agents, such attacks could pose serious threats to physical security, financial security, or overall system integrity.
This paper categorizes the security threats faced by LLM agents into inherited LLM attacks and unique agent-specific threats. The threats inherited from LLMs can be further divided into technical vulnerabilities and intentional malicious attacks. Technical vulnerabilities include issues like hallucinations, catastrophic forgetting, and misunderstandings (Xi et al., 2023), which arise from the initial model creation and are influenced by the model’s structure. These vulnerabilities can cause users to observe incorrect results during prolonged use of LLM agents, affecting user trust and decision-making processes. For instance, AgentBench (Liu et al., [n. d.]) demonstrated that LLMs achieve low success rates between 12% and 14% across real-world tasks, frequently exhibiting hallucination behaviors in knowledge-intensive scenarios, catastrophic forgetting in multi-turn interactions, and misunderstanding in complex reasoning scenarios. On the other hand, intentional malicious attacks deliberately exploit these vulnerabilities to achieve adversarial goals. Representative examples include jailbreaking attacks that bypass alignment safeguards to generate prohibited content, prompt injection attacks that insert malicious instructions into agent inputs to manipulate reasoning processes and outputs, data extraction attacks that induce agents to disclose sensitive training data or model parameters, and inference attacks, which aim to determine whether specific samples were part of the training set (Yao et al., 2023a). Building on these concepts, a variety of concrete attack methods have been developed in recent research and have demonstrated strong attack effectiveness, amplifying the severity of threats and highlighting the complexity and importance of protecting LLM agents.
@@
For the specific threats targeting LLM agents, we are inspired by the workflow of LLM agents, which involves agent perception, thought, and action (Xi et al., 2023). The threats can be categorized into knowledge poisoning, output manipulation and functional manipulation. Knowledge poisoning involves contaminating the training data and knowledge base of the LLM agent, leading to the deliberate incorporation of malicious data by creator. This can easily deceive users with harmful information and even steer them towards malicious behavior. Output manipulation interferes with the content of the agent’s thought and perception stages, influencing the final output. This can cause users to receive biased or deceptive information, crafted to mislead them. Functional manipulation exploits the interfaces and tools used by LLM agents to perform unauthorized actions such as third-party data theft or executing malicious code.
@@
By addressing these aspects, this survey seeks to provide a thorough understanding of the unique challenges posed by LLM agents and contribute to the development of safer and more reliable Artificial General Intelligence (AGI) systems (Zhong et al., 2023).
[…]
Figure 2. Taxonomy of The Emerged Security and Privacy of LLM Agent.
The rest of this paper is structured as follows: Section 2 will delve into the fundamental aspects of LLM agents, including their definition, structure, workflow, and capability. Section 3 will identify and categorizes the emerging threats faced by LLM agents. It discusses both inherited threats from the underlying LLMs and unique agent-specific threats, with detailed examples and scenarios for each category. Section 4 will elaborate on the real-world impacts of the threats. It explores how these threats affect users, environments, and other agents, highlighting the potential consequences of unmitigated risks. Section 5 will review existing mitigation strategies and solutions to address the mentioned threats. Section 6 will discuss gaps in current research and suggests future trends. Section 7 will conclude the article. To provide a clearer visualization of the structure and relationships between the sections, the following Figure 2 presents the framework of this paper.
@@ 2. Foundation of LLM Agent
In this section, we delve into the foundational aspects of LLM agents, exploring their definition, structure, workflow, and capabilities. This exploration is pivotal in understanding the nature of LLM agents.
@@ 2.2. Structure of LLM Agent
  • Personality is a component that defines the tone, style, and interaction manner of an LLM agent. For instance, a tour guide or customer service agent needs to adopt a specific role and perform dialogue tasks in an appropriate manner. In the task of exploring human communities through LLM agent-based societies, agents also need to be endowed with distinct personality traits such as being outgoing, polite, or knowledgeable. Personality assists in simulating realistic emotional expressions and behavioral logic, thereby enabling agents to interact with users and perform tasks consistently and uniquely (Park et al., [n. d.]).
@@
  • Knowledge is the database of information utilized by the LLM agent. It extends the content embedded in the model’s parameters and can include commonsense knowledge, specialized knowledge, and other forms of information, enhancing the agent’s understanding and discussion capabilities in specific tasks (Zou et al., 2025).
  • Memory enables the LLM agent to store and recall information from past interactions. This capability is particularly beneficial in future tasks, helping to retain context and ensure consistency and continuity in interactions, thereby enhancing the overall effectiveness of LLM agents in various applications  (Zhong et al., 2024).
2.3. Workflow of LLM Agent
The fundamental workflow of an LLM Agent consists of a cyclic process comprising Perception, Thought, and Action (Xi et al., 2023). During the Perception phase, the Agent collects and processes information, converting raw data from environmental states and user inputs into Agent-readable formats. In the Thought phase, the Agent analyzes the perceived information and devises potential solutions and corresponding action plans. The Action phase executes specific operations, which may include generating interactive responses, invoking external tools, or performing other relevant tasks. Once these actions are completed, the Agent observes the results through a renewed Perception phase, followed by another round of Thought and Action, as illustrated in Figure 4.
Figure 4. The Fundamental Workflow of LLM Agent:
This Perception–Thought–Action cycle enables the Agent to handle problems dynamically with a human-like ability. Based on this theoretical concept, various implementation frameworks for LLM Agents have emerged. For example, ReAct (Yao et al., [n. d.]) systematically implements this workflow by explicitly demonstrating reasoning (Thought) and tool usage (Action) in dialogues. LangChain (Chase, 2022) engineers this approach through modular design, providing standard components such as Memory and Tools to manage multi-turn dialogues and tool invocations, enabling Agents to effectively complete the Perception-Thought-Action process. Auto-GPT (Significant Gravitas, [n. d.]) and BabyAGI (Nakajima, [n. d.]) further enhance Agent autonomy by introducing task planning and goal decomposition capabilities, enabling Agents to continuously perceive results, think through next steps, and plan after action execution, forming a complete autonomous cycle until goal completion. Platforms like AgentGPT (Python, [n. d.]) and MetaGPT (Hong et al., 2023) focus on improving user interaction and visualization, allowing multiple Agents to collaborate on more complex tasks, thereby extending the Perception–Thought–Action loop to a broader, more sophisticated scale.
@@ 2.4. Capability of LLM Agent
  • Integration with Other AI Systems. LLM agents can also be integrated with different AI systems, like image generators, to offer a more comprehensive set of capabilities, demonstrating their versatility in various applications (Wu et al., 2023).
2.5. Case Study on the Structure, Workflow and Capability of LLM Agent
@@
Figure 5. Simulation Environment and LLM Agent Components:
To better understand the structure, workflow, and capabilities of LLM agents, we employ the town scenario composed of LLM agents, as proposed by  (Lin et al., 2023), for a more detailed introduction. Within this virtual town, each LLM agent consists of multiple key components that define its behavior and operational capabilities. As shown in Figure 5(b), Eva, a store employee agent, integrates structured elements that allow her to autonomously manage store operations and interact with users.
At her core, the LLM Engine utilizes models such as GPT-3.5-turbo and GPT-4, and the project described in (Lin et al., 2023) supports integrating customized models tailored to specific tasks. This enables natural language understanding and response generation, allowing Eva to process customer inquiries and execute intelligent actions smoothly. Her instructions define standardized interaction patterns and response structures, ensuring consistency in her communication. For example, Eva may be instructed to always greet customers with “Hello! How can I assist you today?” before answering their inquiries or to structure product availability responses in a JSON format like {‘‘product’’: ‘‘item_name’’, ‘‘status’’: ‘‘in_stock’’} for seamless integration with the store’s system. Her personality, defined as cheerful, friendly, patient, and efficient, shapes how she engages with customers, ensuring a positive interaction experience. To facilitate these interactions, Eva utilizes the interface provided by the virtual town, a pixelated visual map where users control an agent to navigate the environment and engage in dialogue-based exchanges, creating an intuitive and immersive simulation experience. Her knowledge base enables her to retrieve structured product details such as ingredients, shelf life, and stock levels, allowing her to provide accurate and context-aware responses. Eva’s memory allows her to retain past customer interactions and preferences, enabling her to offer personalized recommendations and improve service quality. Additionally, she is equipped with tools such as a calculator for transaction processing and a scroll viewer for inventory tracking, optimizing her efficiency in managing store operations.
These integrated components work together within Eva’s workflow, allowing her to perceive customer needs, analyze stock data, and execute appropriate actions autonomously like restocking or recommending promotions, continuously refining her behavior through feedback loops.
LLM agents across the virtual town assume diverse roles, demonstrating impressive autonomy and task-handling capabilities. Eva handles customer inquiries, manages inventory through APIs, and processes orders autonomously. She provides personalized recommendations, generates promotional content, adjusts prices, and manages returns, while escalating complex cases to human managers when necessary. By integrating these functions, her role extends beyond physical store operations to online order processing, demonstrating her versatility and integration capabilities.
Through this structured design, Eva enhances store efficiency, optimizes inventory management, and provides an improved shopping experience within the virtual town.
@@ 3. Sources of Threats for LLM Agents
As LLM agents increasingly permeate various industries, serving roles from knowledge query tools to being integrated within robots for aiding in daily human activities, these advanced AI systems have brought unprecedented convenience and benefits to users. However, the widespread adoption and multifunctional capabilities of LLM agents, while offering significant advantages, have also exposed vulnerabilities in their security and reliability. The extensive data resources and potential economic value covered by these systems have rendered them a target for illicit exploitation by malevolent entities. As illustrated in Figure  6, the diagram depicts the potential sources of threats for LLM agents.
It is crucial to understand the sources and nature of these threats because they not only directly impact the security of LLM agents, but may also indirectly affect broader aspects, including the privacy and security of humans, the environment, and other agents. In subsequent sections, we will explore in detail the impacts of these threats and discuss measures that can be taken to mitigate these effects, thereby protecting individuals, the environment, and other agents from potential harm.
Figure 6. The Sources of Threats for LLM Agents:
@@ 3.1. Inherited Threats from LLM
To better illustrate the impact pathways of these threats, we align them with the internal components of LLM agents introduced in Section 2.2 (e.g., LLM Engine, Instruction, Interface, Knowledge, and Memory). This mapping clarifies how technical vulnerabilities and malicious behaviors emerge and propagate from the foundational architecture of LLM agents.
These two types of threats are distinct yet interconnected. Technical vulnerabilities arise from technical limitations during model development rather than malicious intent. Conversely, malicious attacks are intentional actions by external entities aimed at exploiting these vulnerabilities to compromise LLM agents. Despite their different origins and motives, technical vulnerabilities provide opportunities for attackers, enabling them to develop more sophisticated strategies and exposing LLM agents to various security and privacy risks.